OpenStudy (anonymous):

Solve x2 + 4x − 12 = 0. x = 2, x = −6 x = −2, x = −6 x = 3, x = −4 x = −3, x = 4

Mehek (mehek14):

what do you think?

OpenStudy (anonymous):

Do you need help solving it?

OpenStudy (anonymous):

Yes >_>

OpenStudy (anonymous):

Do you know how to start?

Mehek (mehek14):

first factor x2 + 4x − 12

OpenStudy (anonymous):

Have you learned the box method?

OpenStudy (anonymous):

I think its B but I could be wrong or right ;^)

Mehek (mehek14):

B is wrong

Mehek (mehek14):

close but wrong

OpenStudy (anonymous):

In that case it must be A since its the second closest

Mehek (mehek14):

when you factor x2 + 4x -12 you get (x + 6) • (x - 2) so x + 6 = 0 and x - 2 = 0 solve for x

OpenStudy (anonymous):

Lol are you just going to guess all of them?

Mehek (mehek14):

solve for x when x + 6 = 0 x - 2 = 0

OpenStudy (anonymous):

No it wasn't my plan @200205650

OpenStudy (anonymous):

Alright well mehek helped with the factoring step. so now you just need to solve for x.

OpenStudy (anonymous):

Once you find that then you have your answer! :)

OpenStudy (anonymous):

Ok thank you all :D

Mehek (mehek14):

anytime ^_^

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

My second guess was right btw it was A lol

OpenStudy (anonymous):

Lol yes it was.